# How Do Multilingual Models Remember? Investigating Multilingual Factual Recall Mechanisms

### v1

@@ 4 Factual Recall Components
Like Geva et al. (2023), we use attention knockout to visualize the propagation of information to the last token. Based on insights from the causal analysis, we intervene in the last token self-attention in XGLM and cross-attention for mT5. Specifically, we set the attention score to be zero between the last token and another set of tokens $\{t\}$. Given that the contributions of attention blocks are incremental, we perform this intervention within a window of size $w$ around layer $l$. The set of tokens $\{t\}$ consists of (1) subject tokens, (2) non-subject tokens (relation tokens), and (3) the last token (in mT5, the last token refers to the last token in the decoder input). We use a window of size 6 for XGLM and of size 4 for mT5, following Geva et al. (2023) to knock out around 18% of layers. We then measure the change in probability of the originally predicted first token (without knockout), aggregating the results across examples by calculating the relative difference (i.e., the change in probability proportional to the original probability). Thus, a big drop in relative probability indicates a critical flow of information at that layer, from the selected tokens to the last token.
We present the results in Figure 1 (plots per language in Appendix C). For XGLM model, we find broadly similar dynamics to those of GPT (Geva et al., 2023) with a critical flow from the subject at later layers around 25, and negligible flow from the last token to itself. However, there is a strong propagation of information from the non-subject tokens throughout the early and middle layers, shown by the drop of -80% in probability, which reaches -100% for some languages in layers 5-10.888GPT2-xl attention to non-subject tokens is more pronounced in the early layers (-40%), decreasing from there. GPT-J attends to non-subject tokens through early and mid layers but with a weaker drop (-20%) (Geva et al., 2023). Most languages follow a similar pattern, but ru benefits from the knockout to the subject token in early layers; for ar and ko, XGLM attends strongly to non-subject tokens until the end. The mT5 results show a different behavior. First, the last token does attend to itself, a pattern that is more pronounced in the early and later layers, but less so around the middle layers. [^9] We hypothesize that this is probably due to the last token encoding information about which sentinel token is generated at this step (indicating where in the input it should fill), since we see an almost identical curve for all the languages when the last token cannot attend to itself or when it cannot attend to the sentinel token in the decoder (Figure 14). Second, the last token attends to non-subject tokens across all layers, with a weaker drop at the early layers compared to XGLM. Finally, similar to XGLM, there is critical information flowing from the subject token in the later layers. However, this drop is larger and occurs even later in the computation compared to XGLM. Most languages behave similarly, with the exception of ja and ko, exhibiting less flow from last and non-subject tokens in the early layers.
Figure 2: Extraction rates. Left XGLM, right mT5.
@@
Like Geva et al. (2023), we measure extraction events at each layer of the model. Using the notation from §3, let $h^{l}$ be the representation of the last token at layer $l$, and let $E$ be the embedding matrix, then the predicted token is $o^{*}=\arg\,\max(Eh^{L})$. We say there is an extraction event at the self-attention at layer $l$ if $\arg\,\max(Es^{l})=o^{*}$ (similarly for $c^{l}$ and $f^{l}$). The extraction rate is the proportion of examples for which an extraction event occurs at a given layer.
Figure 2 shows the extraction rates for both models, with a detailed language breakdown in Appendix D. A key difference between XGLM and GPT-2 (Geva et al., 2023) is the prominence of MLP modules in the extraction of the object in XGLM. We note that this is not simply because the MLP is forwarding the already extracted object from the preceding attention layer. We plot in Figure 17 how often an MLP has an extraction event while the preceding attention has not, and we find a similar plot to the one in Figure 2, where the MLP has a higher peak in layer 25 than the attention. Additionally, we observe that for English (Figure 15), in line with the results for GPT-2, the attentions have higher extraction rate than the MLPs.
For mT5, the cross-attention layers are the main mechanism for extraction, although the MLP layers also contribute to a lesser degree, resembling GPT-2’s behavior. However, for certain languages (he, fa, ko, and ru), the MLPs show comparable or even higher extraction rates (Figure 16). Finally, we observe that in mT5, extraction events occur later in the computation, with peaks at layers 19 and 22.
5 Information Flow Through Patching
Finally, we study the role of language in the recall prediction of transformer models. We employ activation patching (Wang et al., 2022) by intervening in the model’s computation to identify critical steps. Through these analyses, we contribute three key elements to the understanding of the recall process in transformer models: (1) we localize exactly at which layer the relation information flows to the last token; (2) we identify the extraction mechanism in multilingual LLMs; and (3) we explore how language plays a role in recall prediction.
Figure 3: Patching strategy. The patch example (left) and the context example (right). The patch of this transformer is performed at layer 0 and it affects the computation of layer 1, thus the output is perturbed.
@@ 5 Information Flow Through Patching
Let $h^{l}$ be the representation at layer $l$ of the last token in the input. For a given input $p$, we take $h^{l}_{p}$ and patch it into the forward pass of another input $c$ at the same layer; that is, we set $h^{l}_{t,c}\leftarrow h^{l}_{t,p}$, and then continue the forward pass. We refer to the example $p$ as the patch example, and $c$ as the context example (See Figure 3). By patching the last token representation in each layer and measuring the model’s predicted output, we can study when the representation contains the information about relation, language, and predicted object, and how cross-lingual these representations are.
@@
An input example expresses a relation $r$ and a subject $s$ in language $\mathcal{L}$. We conduct three experiments where the patch and context examples share certain input characteristics. In particular, we test three different conditions: (1) $\{=\mathcal{L},\neq r,\neq s\}$, where the language is the same but the relation and subject differ; (2) $\{\neq\mathcal{L},=r,\neq s\}$, where the relation remains the same but the language and subject differ; and (3) $\{\neq\mathcal{L},\neq r,=s\}$, where the subject is the same but the language and relation differ. (1) allows us to study how the relation is encoded while controlling for the language, whereas (2)-(3) help us explore how the interaction between relation and language affects the extraction of the object from the subject. In all experiments, English is used as the patch language, $\mathcal{L}_{p}=\text{en}$.
@@
Let the input of the patch be $(\mathcal{L}_{p}(r_{p}),\mathcal{L}_{p}(s_{p}))$ and the input of the context be $(\mathcal{L}_{c}(r_{c}),\mathcal{L}_{c}(s_{c}))$. When there is no intervention, the model correctly predicts $\mathcal{L}_{p}(o_{r_{p},s_{p}})$ and $\mathcal{L}_{c}(o_{r_{c},s_{c}})$ respectively. [^10] For simplicity, we may refer to the object as $\mathcal{L}_{p}(o_{p})$ or $\mathcal{L}_{c}(o_{c})$ when $r$ and $s$ are from the same input. To analyze the effect of patching, we study both the output probabilities and the final predictions. First, we evaluate the probability of the originally predicted token, $\mathcal{L}_{p}(o_{p})$ and $\mathcal{L}_{c}(o_{c})$, and similar to §4 we aggregate across examples by calculating the relative difference (the change in probability proportional to the original (unpatched) probability). Second, we examine the predicted token, and we check whether it corresponds to $\mathcal{L}_{p}(o_{c})$, the context answer but in the patch language, or to $\mathcal{L}_{c}(o_{p})$, the patch answer but in the context language. Note that we consider only pairs of patch-context examples for which we have the corresponding objects, and which do not share the same spelling. [^11] We limit ourselves to pairs where: (1) mParaRel has the expected object in the other language (e.g. that we have the aliases of the context object in the patch language $\mathcal{L}_{p}(o_{c})$); and (2) the prediction can only match one of the object aliases sets, e.g. to say that the predicted token $t=\mathcal{L}_{p}(o_{c})$ then $t\neq\mathcal{L}_{c}(o_{c})$, which may happen if the languages share the spelling of the object $o_{c}$. For example, “Asia” is spelled the same in English and Spanish. As a result, the number of examples varies when we study output probabilities, predictions of $\mathcal{L}_{c}(o_{p})$, or predictions of $\mathcal{L}_{p}(o_{c})$. See Table 2 for the number of examples for each model and experiment.
@@
First, we analyze the pairs of examples with $\{=\mathcal{L},\neq r,\neq s\}$. [^12] For example, $r_{p},s_{p}=$“The capital of France is” and $r_{c},s_{c}=$“The language spoken in Germany is”. Then, $o_{r_{p},s_{c}}=$ Berlin and $o_{r_{c},s_{p}}=$French. We study $\sim$1000 examples for which mParaRel has the objects $o_{r_{c},s_{p}}$ and $o_{r_{p},s_{c}}$. [^13] The total number of examples that have different spellings when checking for $o_{r_{p},s_{c}}$ is 984 (994) in XGLM (mT5), and for $o_{r_{c},s_{p}}$ is 974 (990). We present the prediction results in Figure 4, with corresponding probability plots in Figure 20. We observe that the relation information is propagated to the last token representation by layer 13 in XGLM and layer 16 in mT5, since the patching at that point starts producing the object corresponding to the relation in the patch $o_{r_{p},s_{c}}$. By layer 22 in XGLM and layer 18 in mT5, we observe that the object has been fully extracted and is encoded within the last token representation, as predictions from these layers correspond to $o_{r_{p},s_{p}}$. Layer 22 in XGLM aligns with the first attention peak in the extraction rate of English (Figure 15), and layer 18 in mT5 matches the first cross-attention peak (Figure 16). This shows that the extraction measured by the vocabulary projection indeed represents the point at which the object is extracted and encoded in the last token representation. If the object was encoded earlier and only decoded with the extraction event, we would see $o_{r_{p},s_{p}}$ being predicted from patches at earlier layers.
Figure 4: Counts of examples where the patch produces the object prediction. Top is XGLM, bottom mT5.
@@
We have localized the layers where the relation representation flows to the last token. Now, we analyze when the language of the input text propagates to the last token. Is the language information entangled to the subject or the relation representation? Or is the language encoded separately from these two? Here, we study pairs of examples with $\{\neq\mathcal{L},=r,\neq s\}$. [^14] For example: “France’s capital is” and “La capital de Alemania es” (Gloss: “The capital of Germany is”). If we obtain $\mathcal{L}_{p}(o_{c})$, this suggests that the language is encoded in the last token representation before the extraction happens. On the other hand, if we obtain $\mathcal{L}_{c}(o_{p})$, we could infer that the language is encoded after the extraction.151515 After filtering examples with similar object spellings, the total number of pairs we have per language is around 150-800.
@@
We present the probability of the context answer token $\mathcal{L}_{c}(o_{c})$ in Figure 5 (the probability of $\mathcal{L}_{p}(o_{p})$ and per language plots are provided in Appendix E.2). First, we observe that the probability of $\mathcal{L}_{p}(o_{p})$ increases sharply around layers 21-25 for XGLM and around 18-21 for mT5. For XGLM, regarding the probability of $\mathcal{L}_{c}(o_{c})$, we observe that patching in the early layers generally hurts the model’s performance for most languages. However, patching in the middle layers (10-20) either increases the probability or results in only a minimal decrease. Given that the relation is the same in both examples, this suggests that by these middle layers, $r$ is encoded in $h^{l}_{p}$, and in the subsequent layers, the subject and language of the context are integrated to yield the final prediction, $\mathcal{L}_{p}(r)+\mathcal{L}_{c}+s_{c}=\mathcal{L}_{c}(o_{c})$. Moreover, for some languages the relation representation from the patch is better than the one constructed using the context input, as the relative probability is positive.
Figure 5: Relative probability change of $\mathcal{L}_{c}(o_{c})$ when patching the last token representation between pairs of examples with the same relation and a different subject (Top XGLM, bottom mT5).
In the case of mT5, however, the probability of $\mathcal{L}_{c}(o_{c})$ consistently decreases, plateauing in the middle layers before dropping to zero by layer 22. From the attention knockout analysis, we recall that in mT5, the subject is integrated into the last token representation only after layer 15, while the relation is consistently represented throughout the middle layers. Therefore, these patching results imply that the relation encoded in the last token from the patch input does not help in retrieving the correct context object in the target language in mT5, whereas it proves useful in XGLM.
Figure 6: Histogram of patches that cause the context answer to be predicted in the patch language, $\mathcal{L}_{p}(o_{c})$ (Top XGLM, bottom mT5).
@@
In terms of predicted tokens, we find that across-the-board $\mathcal{L}_{c}(o_{p})$ is rarely predicted while $\mathcal{L}_{p}(o_{c})$ appears frequently (Table 3). This suggests that the last token representation encodes the patch language $\mathcal{L}_{p}$ but not yet the object, as the object is derived from the subject context. Figure 6 shows the layers where $\mathcal{L}_{p}(o_{c})$ is predicted (plot for $\mathcal{L}_{p}(o_{c})$ in Figure 31). As discussed in §4, the extraction event typically happens around layers 18-28 in XGLM and 16-22 in mT5. Therefore, we observe that the language is attended to at the beginning of the extraction phase, because if we patch earlier, the output is in the context language. Then, the language is not attended anymore but rather the language information that already flowed to the last token is the one used for the rest of the extraction phase. We can interpret the last token representation as containing the “task” (Todd et al., 2024), the relation that needs to be extracted and in which language, which is used in the extraction event. This task vector can be transferred to contexts in another language in both models, as the task vector is constructed from the patch input and used in the context subject representation to predict $\mathcal{L}_{p}(o_{c})$.
@@
We just saw that the representation from the patch will encode at some point the output language but not yet the object. In the previous experiment both examples had the same relation, so it could be that we observe the prediction $\mathcal{L}_{p}(o_{r,s_{c}})$ because the relation is language specific and encodes the output language. To analyze if this is the case, we now apply patching on examples with different languages and relations but the same subject $\{\neq\mathcal{L},\neq r,=s\}$. [^16] For example: “[Y]’s capital is” and “El idioma oficial de [Y] es” (Gloss: “The official language of [Y] is”). The number of examples is around 150-300 per language, while for XGLM in uk and tr is around 30-70 and only 6 in ko (Table 2). When analyzing the predictions of $\mathcal{L}_{c}(o_{p})$ and $\mathcal{L}_{p}(o_{c})$, as we restrict the examples to have different spellings, the total number of pairs is around 75-250 depending on the language.
Figure 7: Relative probability change of $\mathcal{L}_{c}(o_{c})$ when patching the last token representation between pairs of examples with the different relation and the same subject (Top XGLM, bottom mT5).
@@
We present the probability of the original token $\mathcal{L}_{c}(o_{c})$ and $\mathcal{L}_{p}(o_{p})$ in Figures  7 and 26, respectively. We observe that the probability of $\mathcal{L}_{p}(o_{p})$ starts increasing in layers 10-15 for es, vi, tr, and ko while the increase starts from layers 15-20 for ru, uk, and ar. In mT5, this increase is more uniform across languages and occurs slightly later. As for $\mathcal{L}_{c}(o_{c})$, unlike the previous experiment, we observe that in XGLM, the probability decreases early for all languages (except tr), plateaus around the middle layers, and then drops to zero by the mid-layer range. In mT5, similar to the previous experiment, the probability of $\mathcal{L}_{c}(o_{c})$ drops at the beginning but, instead of plateauing, it continues to decline until it reaches zero around layer 17.
@@
In terms of predictions, we find that $\mathcal{L}_{c}(o_{p})$ is frequently predicted for XGLM (Figure 8), while $\mathcal{L}_{p}(o_{c})$ appears but less often (see Figure 31 and Table 3). This suggests that in layers 15-21, the patched last token representation primarily captures the relation $r_{p}$, without yet encoding the output language or subject information. Therefore, in XGLM, the relation and language representations are disentangled, as the relation flows to the last token before the output language does. Allowing the relation to be combined with different languages.
@@
As for mT5, we observe very few examples where $\mathcal{L}_{c}(o_{p})$ or $\mathcal{L}_{p}(o_{c})$ are predicted, which aligns with the findings of the two former experiments, where we see that the relation is encoded in the last token in layers 15-21 (Figure 4), and the language flows to the last token around layer 15-19 (Figure 6). We conclude that both the relation and language flow to the last token around the same time, and thus, in this experiment, we cannot see a disentangled behavior. This presents an interesting contrast with XGLM. The decoder in mT5 has access to the same encoder representations throughout all its layers, so it does not need (and thus does not learn) to attend to these earlier or in different stages. By contrast, in a decoder-only model like XGLM, the last token has access to representations that evolve across layers, so it learns to attend to relevant information when it becomes most salient. Nonetheless, we cannot reach a definite conclusion on whether the language representation in mT5 is entangled or not to the relation representation. More detailed analysis of what is being attended when the relation flows and when the language flows should be performed in future work.
@@
Figure 8: Histogram of patches that cause the prediction to be $\mathcal{L}_{c}(o_{p})$.

### v2

@@ 4 Factual Recall Components
[…]
Figure 1 presents the results with plots per language in Appendix D. Firstly, the results show that in each model the information flows fairly similarly for all the languages, and similar patterns to those found in English GPT are present (Geva et al., 2023). On the one hand, the subject information flows to the last token most critically at the later layers in all models. On the other hand, the relation flows throughout all the layers, with EuroLLM and mT5 having more similar curves. Nevertheless, we note two differences from the conclusions reached with English GPT (Geva et al., 2023): (1) the propagation of non-subject tokens to the last token does not strictly precedes the subject propagation, in XGLM the drop is the highest in the earlier layers but it continues to be important until the end, and in mT5 and EuroLLM this information has two peaks, in the early and later layers; and (2) the last token of mT5 encodes critical information that flows from one layer to the next through the cross-attention (as opposed to the negligible flow of the last token in the decoder-only models). [^10] We hypothesize this may occur because the last token encodes which sentinel token is being generated, indicating where to fill in the input. Since we see an almost identical curve when the last token cannot attend to the sentinel token in the decoder (Figure 24).
Figure 2: Extraction rates. Left XGLM, right EuroLLM, bottom mT5.
@@
Following Geva et al. (2023), we measure extraction events at each layer. Let $h^{l}$ be the representation of the last token at layer $l$, and let $E$ be the embedding matrix; the predicted token is $o=\arg\,\max(Eh^{L})$. An extraction event occurs at layer $l$ if $\arg\,\max(Es^{l})=o$ (similarly for $c^{l}$ and $f^{l}$). The extraction rate is the proportion of examples for which an extraction event occurs at a given layer. Figure 2 shows the extraction rates for each model, with a detailed language breakdown in Appendix E.
Our results demonstrate that extraction events can be detected in multilingual LLMs, though rates vary across languages. A key finding is the prominence of MLP modules in object extraction for multilingual decoder-only models (XGLM and EuroLLM). To rule out the possibility that MLPs simply forward extracted objects from preceding attention layers, we measure how often MLPs perform an extraction without prior attention extraction (Figure 28). We find that MLPs indeed perform the extraction for most languages, though in English, attention modules dominate, aligning with GPT-2 results (Geva et al., 2023). In mT5, cross-attention layers drive the extraction, with MLPs playing a secondary role, resembling GPT-2’s behavior. Additionally, in mT5 the extraction events occur in later layers, with peaks in 19 and 22.
The results show that in multilingual LLMs, information flow from non-subject tokens to the final token is less localized compared to English LLMs, possibly because the model processes relational and language information at different stages. Additionally, the extraction mechanism in multilingual LLMs is more complex, involving both attention and MLP modules. This implies that editing techniques should focus not only on early MLPs enhancing subject representations but also on the later MLPs that extract the object.
5 Language Information Flow
Previous analyses do not provide insights into how the language is included and used during recall. In this section, we use activation patching, similarly to Dumas et al. (2024), to: (1) disentangle when the relation information flows to the last token and when the language information flows, (2) identify which tokens contribute the language information, and (3) interpret the last token representation as a Function Vector Todd et al. (2024) [^11] Function Vectors (FV) were originally defined as the mean vector of outputs from specific attention heads. Here, we interpret the last token representation as an FV because it triggers a specific execution for different contexts..
Figure 3: Patching strategy. The patch example’s last token is inserted in the context example’s forward pass, which perturbs the output of subsequent layers.
(a) Counts of examples where the patch produces the object prediction.
(b) Histogram of patches that cause the context answer to be predicted in the patch language, $\mathcal{L}_{p}(o_{c})$.
Figure 5: Prediction tokens from (a) $\{=\mathcal{L},\neq r,\neq s\}$ and (b) $\{\neq\mathcal{L},=r,\neq s\}$ for XGLM, EuroLLM, mT5.
@@ 5 Language Information Flow
Let $h^{l}_{t}$ be the representation at layer $l$ of the last token in the input. For a given input $p$, we take $h^{l}_{t,p}$ and patch it into the forward pass of another input $c$ at the same layer; that is, we set $h^{l}_{t,c}\leftarrow h^{l}_{t,p}$, and then continue the forward pass. We refer to the example $p$ as the patch example, and $c$ as the context example (see Figure 3). By patching the last token representation in each layer and measuring the model’s predicted output, we can study when the representation contains the information about relation, language, and predicted object, and how cross-lingual these representations are.
@@
An input example expresses a relation $r$ and a subject $s$ in language $\mathcal{L}$, which we note $(\mathcal{L}(r),\mathcal{L}(s))$. We conduct three experiments where the patch and context examples share certain input characteristics. In particular, we test: (1) $\{=\mathcal{L},\neq r,\neq s\}$, where the language is the same but the relation and subject differ; (2) $\{\neq\mathcal{L},=r,\neq s\}$, where the relation remains the same but the language and subject differ; and (3) $\{\neq\mathcal{L},\neq r,=s\}$, where the subject is the same but the language and relation differ. (1) allows us to study how the relation is encoded while controlling for the language, whereas (2)-(3) help us explore how the interaction between relation and language affects the extraction of the object from the subject. In all experiments, English is used as the patch language, $\mathcal{L}_{p}=\text{en}$.
@@
Let the patch input be $(\mathcal{L}_{p}(r_{p}),\mathcal{L}_{p}(s_{p}))$ and the context input be $(\mathcal{L}_{c}(r_{c}),\mathcal{L}_{c}(s_{c}))$. Without intervention, the model correctly predicts $\mathcal{L}_{p}(o_{r_{p},s_{p}})$ and $\mathcal{L}_{c}(o_{r_{c},s_{c}})$ respectively. [^12] For simplicity, we may refer to the object as $\mathcal{L}_{p}(o_{p})$ or $\mathcal{L}_{c}(o_{c})$ when $r$ and $s$ are from the same input. To analyze patching effects, we examine both output probabilities and predicted tokens. For probabilities, we aggregate across examples by calculating the change relative to the original (unpatched) probability, as in §4. For predicted tokens, we check if they match the patch or context object, or a variant with swapped language or relation (e.g., if $\mathcal{L}_{p}(o_{c})$ is predicted in the setups (2)-(3)). We only consider valid patch-context pairs where $o_{r_{p},s_{c}}$ and $o_{r_{c},s_{p}}$ exist, and for predicted token analysis, we enforce distinct spellings. [^13] E.g. to claim the predicted token is $t=\mathcal{L}_{p}(o_{c})$ we require $t\neq\mathcal{L}_{c}(o_{c})$, as languages may share spellings (e.g., “Asia” in English and Spanish). Consequently, the number of examples varies across analyses of output probabilities and token predictions. See Table 4 and 6 for the number of examples for each model and experiment.
Figure 6: Relative probability change of $\mathcal{L}_{c}(o_{c})$ when patching $\{\neq\mathcal{L},=r,\neq s\}$, for XGLM, EuroLLM, mT5.
Figure 7: Relative probability change of $\mathcal{L}_{c}(o_{c})$ when patching $\{\neq\mathcal{L},\neq r,=s\}$, for XGLM, EuroLLM, mT5.
@@
First, we analyze the pairs of examples with $\{=\mathcal{L},\neq r,\neq s\}$. For example, $r_{p},s_{p}=$“The capital of France is” and $r_{c},s_{c}=$“The language spoken in Germany is”. Then, $o_{r_{p},s_{c}}=$ Berlin and $o_{r_{c},s_{p}}=$ French. We sample 1000 examples for which mParaRel has the objects $o_{r_{c},s_{p}}$ and $o_{r_{p},s_{c}}$.
We present the prediction results in Figure 5(a), and probability plots in Figure 32. The relation information is contained in the last token representation when the green curve increases, since patching at that point starts producing the object corresponding to the relation in the patch $o_{r_{p},s_{c}}$. When the green curve decreases and the blue curve increases, the object has been fully extracted and is encoded in the last token vector, as predictions from these layers correspond to $o_{r_{p},s_{p}}$ and no information is taken from the context $c$. The layers where this happens align with the peaks of the extraction rate in English (Figure 25-27). This shows that the extraction measured by the vocabulary projection represents the point at which the object is extracted and encoded in the last token representation. If the object was encoded earlier and only decoded with the extraction event,$o_{r_{p},s_{p}}$ would be predicted from patches at earlier layers.
@@
We have localized the layers where the relation representation flows to the last token. Now, we analyze when the language of the input text propagates to the last token. Is the language information entangled to the subject or the relation representation? Here, we study pairs of patch-context with $\{\neq\mathcal{L},=r,\neq s\}$, e.g., “France’s capital is” and “La capital de Alemania es” (Gloss: “The capital of Germany is”). If we obtain $\mathcal{L}_{p}(o_{c})$ (“Berlin”), it would suggest that the language is encoded in the last token representation before the extraction happens. On the other hand, if we obtain $\mathcal{L}_{c}(o_{p})$ (“París”), we could infer that the language is encoded after the extraction.
@@
We present the probability of the context answer token $\mathcal{L}_{c}(o_{c})$ in Figure 7.141414The probability of $\mathcal{L}_{p}(o_{p})$ and per language plots are provided in Appendix F.2). We observe that for decoder-only models, patching in the early layers generally hurts the model’s performance for most languages, while, patching in the middle layers either increases the probability or results in only a minimal decrease. Given that the relation is the same in both examples, this suggests that by these middle layers, $r$ is encoded in $h^{l}_{p}$, and in the subsequent layers the subject and language of the context are integrated to yield the final prediction, $\mathcal{L}_{p}(r)+\mathcal{L}_{c}+s_{c}=\mathcal{L}_{c}(o_{c})$. Moreover, for some languages the relation representation from the patch is better than the one constructed using the context input, as the relative probability is positive. In the case of mT5, however, the probability of $\mathcal{L}_{c}(o_{c})$ consistently decreases, plateauing in the middle layers before dropping to zero. From the attention knockout analysis, we recall that in mT5, the subject is integrated into the last token representation only after layer 15, while the relation is consistently represented throughout the middle layers. Therefore, these patching results imply that the relation encoded in the last token from the patch input does not help in retrieving the correct context object in the context language in mT5, whereas it proves useful in XGLM and EuroLLM.
@@
In terms of predicted tokens, we find across-the-board that $\mathcal{L}_{p}(o_{c})$ is frequently predicted while $\mathcal{L}_{c}(o_{p})$ is not (Table 5). This suggests that the last token representation encodes the patch language $\mathcal{L}_{p}$ but not yet the object, as the object is derived from the context subject. We plot the layers where $\mathcal{L}_{p}(o_{c})$ is predicted in Figure 5(b) ($\mathcal{L}_{c}(o_{p})$ in Figure 47). We can conclude that the language information flows to the last token right before the peak of $\mathcal{L}_{p}(o_{c})$ predictions, because if we patch earlier, the output is in the context language (see the probabilities of $\mathcal{L}_{c}(o_{c})$ in Figure 7). Moreover, these peaks match the beginning of their corresponding extraction phases, thus the language information flows right before the extraction phase.
As a result, we interpret the last token representation as containing a Function Vector (FV), the relation that needs to be extracted and in which language, which is used in the extraction event. The FV can be transferred to contexts in another language, as the FV is constructed from the patch input and is used in the context subject representation to predict $\mathcal{L}_{p}(o_{c})$.
@@
We just saw that for all models the representation from the patch will encode at some point the output language but not yet the object. It could be that we observed the prediction $\mathcal{L}_{p}(o_{c})$ because the relation is language specific and encodes the output language. To analyze if this is the case, we now apply patching on examples with different languages and relations but the same subject $\{\neq\mathcal{L},\neq r,=s\}$, e.g., “France’s capital is” and “El idioma oficial de Francia es” (Gloss: “The official language of France is”).
@@
We observe that the probability of $\mathcal{L}_{c}(o_{c})$ (Figures  7) in decoder-only models, unlike the previous experiment, decreases early for all languages (except tr and ko), plateaus around the middle layers, and then drops to zero by the mid-layer range. For mT5, the probability drops at the beginning but, instead of plateauing as before, it continues to decline until it reaches zero. When compared to the former experiment (Figure 7), this suggests that the relation information is encoded in the middle layers, as $\mathcal{L}_{c}(o_{c})$ decreases earlier when the patch and context have different relations.
@@
In terms of predictions, we find that $\mathcal{L}_{c}(o_{p})$ is frequently predicted for XGLM and EuroLLM (Figure 8), while $\mathcal{L}_{p}(o_{c})$ appears but less often (Table 7). In line with the previous observation, the plot shows that the last token representation in the middle layers where $\mathcal{L}_{c}(o_{p})$ is predicted, primarily captures the relation from the patch $r_{p}$, without yet encoding the output language or subject information (as these are taken from the context). Therefore, in decoder-only models, the relation and language representations are disentangled, as the relation flows to the last token before the output language does. Allowing the relation to be combined with different languages.
@@
As for mT5, we observe very few examples where $\mathcal{L}_{c}(o_{p})$ or $\mathcal{L}_{p}(o_{c})$ are predicted, which aligns with the findings of the two former experiments, where we see that the relation is encoded in the last token in layers 15-21 (Figure 5(a)), and the language flows to the last token around layer 15-19 (Figure 5(b)). We conclude that both the relation and language flow to the last token around the same time, and thus, in this experiment, we cannot see a disentangled behavior. This presents an interesting contrast with decoder-only models. The decoder in mT5 has access to the same encoder representations throughout all its layers, so it does not need (and thus does not learn) to attend to these earlier or in different stages. By contrast, in a decoder-only model, the last token has access to representations that evolve across layers, so it learns to attend to relevant information when it becomes most salient. Nonetheless, we cannot reach a definite conclusion on whether the language representation in mT5 is entangled or not to the relation representation. More detailed analysis of what is being attended when the relation flows and when the language flows should be performed in future work.
@@
Figure 8: Patches that cause the prediction to be $\mathcal{L}_{c}(o_{p})$. Top XGLM, bottom EuroLLM.